# Intrinsic Mixed-state Quantum Topological Order

### v2

@@ A.1 Calculation of the coherent information and mapping to the RBIM
In this section, we present detailed calculation of the coherent information $I_{c}=S(\rho_{f})-S(\rho_{Rf})$, and the derivation of the mapping to the RBIM along the Nishimori line ((4)) Nishimori (1981). In the next two subsections, we calculate the two von Neumann entropy $S(\rho_{f})$ and $S(\rho_{Rf})$ respectively, using the replica trick: $S=-\text{Tr}(\rho\log{\rho})=-\lim_{n\rightarrow 1}\frac{\partial}{\partial n}\text{Tr}(\rho^{n})$.
@@ A.1.3 Critical error rate and classical memory from coherent information
where $\Delta F_{d_{x},d_{y}}$ is the excess free energy with the insertion of a non-contractible defect line. For small $p$, the RBIM is in the ferromagnetic (FM) phase, and the excess free energy of a defect line is extensive, $\Delta F_{\{d_{x},d_{y}\}\neq\{0,0\}}\sim O(L)$, which leads to $I_{c}=2\log 2$. On the other hand, when $p$ is above the error threshold $p_{c}\approx 0.109$, the RBIM undergoes a phase transition to a paramagnetic (PM) phase, and $I_{c}<2\log 2$. This is exactly the same as the situation with single-qubit errors. The classical memory of $\rho_{f}$ is also reflected in the fact that $I_{c}\geq 0$.
@@ A.3 Calculation of the entanglement negativity
$\text{Tr}(\rho^{T_{A}}_{f})^{2n}$ $=2^{-2nN}\prod_{s=1}^{n}\sum_{\tilde{g}^{(s)}\in H}\prod_{s=1}^{n}\sum_{g^{(}{s)}\in G}(1-2p)^{\sum_{s=1}^{n}l_{g^{(s)}}+l_{g^{(s)}\tilde{g}^{(s)}}}y_{A}(\prod_{s=1}^{n}\tilde{g}^{(s)})\text{Tr}\left(\prod_{s=1}^{n}\tilde{g}^{(s)}\right)$ (59)
@@
$\text{Tr}\rho_{f}^{2n}=2^{(1-2n)N}\sum_{\tilde{g}^{1},\cdots,\tilde{g}^{n-1}\in G}\sum_{g^{1},\cdots,g^{n}\in G}(1-2p)^{\sum_{s=1}^{n-1}(l_{g^{(s)}}+l_{g^{(s)}\tilde{g}^{(s)}})+l_{\prod_{s=1}^{n-1}g^{(s)}}+l_{\prod_{s=1}^{n-1}g^{(s)}\tilde{g}^{(s)}}}.$ (60)
@@
$O_{\{\tilde{g}\}}=2^{(1-2n)N}\sum_{g^{1},\cdots,g^{n}\in G}(1-2p)^{\sum_{s=1}^{n-1}(l_{g^{(s)}}+l_{g^{(s)}\tilde{g}^{(s)}})+l_{\prod_{s=1}^{n-1}g^{(s)}}+l_{\prod_{s=1}^{n-1}g^{(s)}\tilde{g}^{(s)}}}.$ (61)
@@
$\text{Tr}(\rho_{f}^{T_{A}})^{2n}$ $=\sum_{\tilde{g}^{1}_{f},\cdots\tilde{g}^{n-1}_{f}\in H_{f}}\sum_{\tilde{u}^{1},\cdots,\tilde{u}^{n-1}\in H/H_{f}}O_{\{\tilde{u}\}},$ (62)
$\text{Tr}\rho_{f}^{2n}$ $=\sum_{\tilde{g}^{1}_{f},\cdots\tilde{g}^{n-1}_{f}\in G_{f}}\sum_{\tilde{u}^{1},\cdots,\tilde{u}^{n-1}\in G/G_{f}}O_{\{\tilde{u}\}}$
@@
Bipartition 2, $p=\frac{1}{2}$. For bipartition 2, the calculation for generic $p$ is much more challenging. This is mainly because $G/G_{f}\neq H/H_{f}$, which means (63) cannot be derived from (62) for generic $p$. As a result, the negativity does depend on $p$ for this bipartition. Here we are mainly interested in the phase with intrinsic mixed-state quantum TO for $p>p_{c}$, so we take the maximally decohered limit, $p=\frac{1}{2}$, in which case the calculation can be greatly simplified.
@@ A.5 Robustness under phase errors
In this section, we aim to discuss the robustness of the intrinsic mixed-state quantum topological order under other noises. We derive the phase diagram in Fig. 2 with additional single-qubit phase errors. Concretely, we consider the following mixed state:
@@
Similar to section A, B, we analyze the property of $\tilde{\rho}_{f}$ by calculating the von Neumann entropy $S(\tilde{\rho}_{f})$ and map it to statistical models.
@@
Applying the same strategy as in section A, B, we can map the relative entropy and coherent information to observables in the RBIM. We directly list the results here.
@@ Appendix B Exact solution to the gapless spin liquid phase of the toric code model through fermionization
We show that this model can be solved exactly using the method introduced in Chen et al. (2018). Firstly, we note that the model has an extensive number of local conserved quantities. $[H,W_{p}]=0$ with $W_{p}=A_{v=p-\bm{\delta}}B_{p}$, which follows from the fact that the $e$ anyons and adjacent $m$ anyons are always created or annihilated in pairs, so we can solve the model in each simultaneous eigenspace of $W_{p}$. Secondly, the role of $ZX$ term is to induce pair creation, annihilation and hopping of $f$ anyons, which are fermions. Then on a infinite lattice or a topologically trivial lattice, in each sector $\{W_{p}=w_{p}\}$, the only degrees of freedom are the $f$ anyons, so we expect in each sector the model can be described by a fermion tight-binding model. We assume the fermions are defined on the vertices of the lattice, with the mapping
@@
$\xi_{k}=\sqrt{[2-h_{xz}(\cos k_{x}+\cos k_{y})]^{2}+[h_{xz}(\sin k_{x}+\sin k_{y})]^{2}},$ (93)
and the ground state energy is $E_{g}=-\sum_{k}\xi_{k}$. For $h_{xz}<1$, the spectrum is gapped and the ground energy is nearly degenerate (with an exponentially small splitting) for the 4 types of boundary conditions. This corresponds to the gapped topologically ordered phase of $H$. For $h_{xz}=1$, the gap closes at $k_{x}=k_{y}=0$, and remain closed for $h_{xz}>1$, with linear dispersion at two Dirac points $k_{x}=-k_{y}=\pm\arccos\frac{1}{h_{xz}}$. So for $h_{xz}>1$ the original model lies in a gapless spin liquid phase, reminiscent of the gapless phase of the Kitaev honeycomb model. In this phase the topological degeneracy is lifted by an algebraically small gap, but the ground state remains long-range entangle.
@@ C.2 Entanglement negativity
$(\rho_{f}^{T_{A}})^{2}$ $=(\frac{1}{2^{N+L+1}})^{2}\sum_{g,g^{\prime},w,w^{\prime}}\sum_{g_{2},g_{2}^{\prime},w_{2},w_{2}^{\prime}}D_{g_{A}^{\prime}}D_{g_{B}}\left[|u_{A},w^{\prime}\rangle|u_{B},w\rangle\otimes\rho_{F}\right]$ (103)
$\langle u_{A},w|\langle u_{B},w^{\prime}|D_{g^{\prime}_{B}}D_{g_{A}}D_{g_{2,B}}D_{g^{\prime}_{2,A}}|u_{A},w_{2}^{\prime}\rangle|u_{B},w_{2}\rangle$
$\rho_{F}^{T_{A}}\otimes\langle u_{A},w_{2}|\langle u_{B},w_{2}^{\prime}|,$

### v3

@@ A.1 Calculation of the coherent information and mapping to the RBIM
In this section, we present detailed calculations of the coherent information $I_{c}=S(\rho_{f})-S(\rho_{Rf})$, and the derivation of the mapping to the RBIM along the Nishimori line ((4)) Nishimori (1981). In the next two subsections, we calculate the two von Neumann entropy $S(\rho_{f})$ and $S(\rho_{Rf})$ respectively, using the replica trick: $S=-\text{Tr}(\rho\log{\rho})=-\lim_{n\rightarrow 1}\frac{\partial}{\partial n}\text{Tr}(\rho^{n})$.
@@ A.1.3 Critical error rate and classical memory from coherent information
where $\Delta F_{d_{x},d_{y}}$ is the excess free energy with the insertion of a non-contractible defect line. For small $p$, the RBIM is in the ferromagnetic (FM) phase, and the excess free energy of a defect line is extensive, $\Delta F_{\{d_{x},d_{y}\}\neq\{0,0\}}\sim O(L)$, which leads to $I_{c}=2\log 2$. On the other hand, when $p$ is above the error threshold $p_{c}\approx 0.109$, the RBIM undergoes a phase transition to a paramagnetic (PM) phase, and $I_{c}$ drops to 00 in the thermodynamic limit, which indicates that $\rho_{f}$ only retains a classical memory. This is exactly the same as the situation with single-qubit errors.
@@ A.3 Calculation of the entanglement negativity
$\text{Tr}(\rho^{T_{A}}_{f})^{2n}$ $=2^{-2nN}\prod_{s=1}^{n}\sum_{\tilde{g}^{(s)}\in H}\prod_{s=1}^{n}\sum_{g^{{\color[rgb]{1,0,0}(s)}}\in G}(1-2p)^{\sum_{s=1}^{n}l_{g^{(s)}}+l_{g^{(s)}\tilde{g}^{(s)}}}y_{A}(\prod_{s=1}^{n}\tilde{g}^{(s)})\text{Tr}\left(\prod_{s=1}^{n}\tilde{g}^{(s)}\right)$ (59)
@@
$\text{Tr}\rho_{f}^{2n}=2^{(1-2n)N}\sum_{\tilde{g}^{(1)},\cdots,\tilde{g}^{(n-1)}\in G}\sum_{g^{(1)},\cdots,g^{(n)}\in G}(1-2p)^{\sum_{s=1}^{n-1}(l_{g^{(s)}}+l_{g^{(s)}\tilde{g}^{(s)}})+l_{\prod_{s=1}^{n-1}g^{(s)}}+l_{\prod_{s=1}^{n-1}g^{(s)}\tilde{g}^{(s)}}}.$ (60)
@@
$O_{\{\tilde{g}\}}=2^{(1-2n)N}\sum_{g^{(1)},\cdots,g^{(n)}\in G}(1-2p)^{\sum_{s=1}^{n-1}(l_{g^{(s)}}+l_{g^{(s)}\tilde{g}^{(s)}})+l_{\prod_{s=1}^{n-1}g^{(s)}}+l_{\prod_{s=1}^{n-1}g^{(s)}\tilde{g}^{(s)}}}.$ (61)
@@
$\text{Tr}(\rho_{f}^{T_{A}})^{2n}$ $=\sum_{\tilde{g}^{(1)}_{f},\cdots\tilde{g}^{(n-1)}_{f}\in H_{f}}\sum_{\tilde{u}^{(1)},\cdots,\tilde{u}^{(n-1)}\in H/H_{f}}O_{\{\tilde{u}\}},$ (62)
$\text{Tr}\rho_{f}^{2n}$ $=\sum_{\tilde{g}^{(1)}_{f},\cdots\tilde{g}^{(n-1)}_{f}\in G_{f}}\sum_{\tilde{u}^{(1)},\cdots,\tilde{u}^{(n-1)}\in G/G_{f}}O_{\{\tilde{u}\}}$
@@
Bipartition 2, $p=\frac{1}{2}$. For bipartition 2, the calculation for generic $p$ is much more challenging. This is mainly because $G/G_{f}\neq H/H_{f}$, which means (63) cannot be derived from (62) for generic $p$. As a result, the negativity does depend on $p$ for this bipartition. Here we are mainly interested in the phase with intrinsic mixed-state TO for $p>p_{c}$, so we take the maximally decohered limit, $p=\frac{1}{2}$, in which case the calculation can be greatly simplified.
@@ A.5 Robustness under phase errors
In this section, we aim to discuss the robustness of the intrinsic mixed-state topological order under other noises. We demonstrate how to obtain the phase diagram in Fig. 2 with additional single-qubit phase errors. Concretely, we consider the following mixed state:
@@
Similar to sections A, B, we analyze the property of $\tilde{\rho}_{f}$ by calculating the von Neumann entropy $S(\tilde{\rho}_{f})$ and map it to statistical models.
@@
Applying the same strategy as in sections A, B, we can map the relative entropy and coherent information to observables in the RBIM. We directly list the results here.
@@ Appendix B Exact solution to the gapless spin liquid phase of the toric code model through fermionization
We show that this model can be solved exactly using the method introduced in Chen et al. (2018). Firstly, we note that the model has an extensive number of locally conserved quantities. $[H,W_{p}]=0$ with $W_{p}=A_{v=p-\bm{\delta}}B_{p}$, which follows from the fact that the $e$ anyons and adjacent $m$ anyons are always created or annihilated in pairs, so we can solve the model in each simultaneous eigenspace of $W_{p}$. Secondly, the role of $ZX$ term is to induce pair creation, annihilation and hopping of $f$ anyons, which are fermions. Then on an infinite lattice or a topologically trivial lattice, in each sector $\{W_{p}=w_{p}\}$, the only degrees of freedom are the $f$ anyons, so we expect in each sector the model can be described by a fermion tight-binding model. We assume the fermions are defined on the vertices of the lattice, with the mapping
@@
$\xi_{k}=2\sqrt{[2-h_{xz}(\cos k_{x}+\cos k_{y})]^{2}+[h_{xz}(\sin k_{x}+\sin k_{y})]^{2}},$ (93)
and the ground state energy is $E_{g}=-\sum_{k}\frac{\xi_{k}}{2}$. For $h_{xz}<1$, the spectrum is gapped and the ground energy is nearly degenerate (with an exponentially small splitting) for the 4 types of boundary conditions. This corresponds to the gapped topologically ordered phase of $H$. For $h_{xz}=1$, the gap closes at $k_{x}=k_{y}=0$, and remain closed for $h_{xz}>1$, with linear dispersion at two Dirac points $k_{x}=-k_{y}=\pm\arccos\frac{1}{h_{xz}}$. So for $h_{xz}>1$ the original model lies in a gapless spin liquid phase, reminiscent of the gapless phase of the Kitaev honeycomb model. In this phase the topological degeneracy is lifted by an algebraically small gap, but the ground state remains long-range entangle.
@@ C.2 Entanglement negativity
$(\rho_{f}^{T_{A}})^{2}=$ $(\frac{1}{2^{N+L+1}})^{2}\sum_{g,g^{\prime},w,w^{\prime}}\sum_{g_{2},g_{2}^{\prime},w_{2},w_{2}^{\prime}}\langle u_{A},w|\langle u_{B},w^{\prime}|D_{g^{\prime}_{B}}D_{g_{A}}D_{g_{2,B}}D_{g^{\prime}_{2,A}}|u_{A},w_{2}^{\prime}\rangle|u_{B},w_{2}\rangle$ (103)
$D_{g_{A}^{\prime}}D_{g_{B}}|u_{A},w^{\prime}\rangle|u_{B},w\rangle\langle u_{A},w_{2}|\langle u_{B},w_{2}^{\prime}|D_{g^{\prime}_{2B}}D_{g^{\prime}_{2A}}\otimes(\rho_{F}^{T_{A}})^{2}$